OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (3, 4) and is parallel to the line y = –x + 6?

OpenStudy (anonymous):

hi king, what nincomp was saying is there is two equation of the line there, but they are parallel to each other, means their slope m are equal

OpenStudy (anonymous):

use the slope of this line y = –x + 6, to get the slope of the other line

OpenStudy (anonymous):

to get the slope of that you can compare it to the equation of y=mx+b y = –x + 6 we see that mx=-x so m=-x/x=-1=slope

OpenStudy (anonymous):

haha ok, thats a good start lol, now use m=-1and the the point (3, 4) to the formula y-y1=m(x-x1)

OpenStudy (anonymous):

hi king are you getting it?

OpenStudy (anonymous):

well might as well continue it lol.... using m=-1and the the point (3, 4)=(x,y) to the formula y-y1=m(x-x1) y-4=-1(x-3) ans or y-4=-x+3 y=-x+3+4 y=-x+7 .....ans

OpenStudy (anonymous):

hello equation of the line is y=-x+b (some slope) that passes through the point (3, 4) ; 4=-3+b b=7 the equation is y = -x +7